OpenStudy (anonymous):

Samantha has a shelf that is 95 over 4 inches wide. How many books can Samantha arrange on the shelf if each book is 5 over 4 inches thick? 29, because the number of books is 95 over 4 multiplied by 5 over 4 28, because the number of books is 95 over 4 divided by 4 over 5 19, because the number of books is 95 over 4 divided by 5 over 4 17, because the number of books is 95 over 4 multiplied by 4 over 5

OpenStudy (anonymous):

d?

OpenStudy (campbell_st):

ok so you are dividing by a fraction \[\frac{95}{4} \div \frac{5}{4} \] the rule is flip the divisor and multiply so you have \[\frac{95}{4} \times \frac{4}{5} = \] now just calculate the answer

OpenStudy (anonymous):

380 / 20